A triangle has a 3-inch side, a 4-inch side, and a 5-inch side. The altitude drawn to the 5-inch side cuts this side into segments of what lengths?
Answer by Boreal(15235)

Can draw the altitude to the hypotenuse and have legs x and 5-x, and define x as the longer.
x^2+a^2=16, for one triangle
(5-x)^2+a^2=9
subtracting, one gets x^2-(5-x)^2=7
this works out to x=3.2 and 5-x=1.8 units inches. These can be shown to be consistent with the Pythagorean theorem for both triangles.